A ball is thrown straight downward with an initial velocity of 0.5 m/s down from a height of 4 meters. What is the magnitude and direction of the ball's velocity 0.7 seconds after it is released?


1
Expert's answer
2021-09-12T19:00:10-0400

We can find the ball's velocity 0.7 seconds after it is released from the kinematic equation:


"v=v_0-gt,""v(t=0.7\ s)=-0.5\ \dfrac{m}{s}-9.8\ \dfrac{m}{s^2}\cdot0.7\ s=-7.36\ \dfrac{m}{s}."

The sign minus means that the ball's velocity directed downward.

Answer:

"v(t=0.7\ s)=7.36\ \dfrac{m}{s}", downward.